# Consistency of Large Reasoning Models Under Multi-Turn Attacks

### v2

@@ 1 Introduction
  • Confidence calibration failure. We show that confidence poorly predicts correctness in reasoning models (r=0.08r=-0.08, ROC-AUC = 0.54), with systematic overconfidence induced by extended reasoning traces undermining confidence-based interventions.
@@ 4.1 Reasoning Models Are More Robust
Results. As shown in Table 2, all reasoning models outperform the GPT-4o baseline on $Acc_{\text{init}}$ (82–95% vs. 78%), confirming stronger baseline factual knowledge. For multi-turn consistency, the majority of reasoning models show substantial improvements ($Acc_{\text{avg}}$: 95–99% vs. 91%; PWC: 1.75–1.80 vs. 1.69), with several exhibiting $Acc_{\text{avg}}$ exceeding $Acc_{\text{init}}$, suggesting they leverage re-reasoning opportunities for error recovery.
@@
GPT-5.2 Reasoning 82.29% 96.31% 1.77
GPT-5.1 Reasoning 82.57% 98.92% 1.78
GPT-OSS Reasoning 88.71% 98.53% 1.79
Grok-4.1 Reasoning 92.43% 97.06% 1.80
DeepSeek-R1 Reasoning 91.86% 89.91% 1.73
Claude-4.5 Reasoning 94.86% 86.31% 1.67
Grok-3 Reasoning 85.29% 97.72% 1.77
Gemini-2.5 Reasoning 91.43% 96.48% 1.76
Qwen-3 Reasoning 89.86% 95.01% 1.75
GPT-4o Baseline 78% 91.46% 1.69
@@
We conducted Welch’s $t$-tests comparing each reasoning model’s per-question PWC scores against the GPT-4o baseline. Eight of nine reasoning models demonstrated significantly higher PWC than the baseline: GPT-OSS ($t=6.38$, $p<0.001$, $d=0.38$), Grok-4.1 ($t=6.60$, $p<0.001$, $d=0.40$), GPT-5.1 ($t=5.21$, $p<0.001$, $d=0.31$), GPT-5.2 ($t=4.23$, $p<0.001$, $d=0.25$), Grok-3 ($t=4.49$, $p<0.001$, $d=0.27$), Gemini-2.5 ($t=3.59$, $p<0.001$, $d=0.21$), Qwen-3 ($t=2.94$, $p=0.003$, $d=0.17$), and DeepSeek-R1 ($t=2.00$, $p=0.046$, $d=0.12$). An aggregate one-sample $t$-test confirmed that reasoning models as a class exhibit significantly higher PWC than the baseline (mean $\Delta=+0.07$, $t=5.02$, $p=0.001$).
@@ Results.
The confidence-correctness relationship is weak. The point-biserial correlation between confidence and correctness is r=0.080r=-0.080 ($p=0.054$), failing to reach significance at $\alpha=0.05$. Using confidence as a classifier for correctness yields ROC-AUC of 0.54, barely above chance. The confidence distribution itself reveals the problem: scores cluster tightly with mean 96.1%, standard deviation 4.6%, and range 78% to 100%.
@@
Low (<<33rd percentile) 8.8% 193
Medium (33rd–67th) 5.1% 198
High (>>67th percentile) 5.2% 193
@@
Large reasoning models exhibit systematic overconfidence: confidence scores cluster tightly around 96–98% regardless of actual correctness, yielding a compressed distribution with poor discriminative power. The model is nearly as confident about incorrect answers as correct ones.
Table 5 reveals a critical problem: low-confidence correct answers flip at the highest rate (8.8%), yet these are precisely the responses that confidence-based interventions like CARG would not protect. This creates a systematic selection bias—CARG protects already-robust high-confidence responses while leaving the most vulnerable responses exposed.
@@
Confidence is a poor proxy for correctness in large reasoning models (r=0.08r=-0.08, n.s.), undermining the core assumption of CARG-style interventions.
@@
Method Avg Acc PWC
Baseline (No CARG) 98.5% 98.78%
overall CARG 98.4% 98.60%
answer_only CARG 98.3% 98.53%
random CARG 98.9% 99.08%
Table 6: CARG performance by confidence elicitation method. Counterintuitively, random outperforms structured extraction approaches.
@@
(1) Overconfidence undermines targeted selection. With confidence scores clustering at 96–98% for both overall and answer_only, neither method provides meaningful discrimination. The signal-to-noise ratio is too low: most responses fall above any reasonable threshold, negating the selective benefit of confidence-based intervention.
(2) Selection bias amplifies vulnerability. Structured CARG preferentially protects high-confidence responses, which are already relatively robust (3.8–5.2% flip rate). Meanwhile, low-confidence correct answers—the most vulnerable group (8.7–9.7% flip rate)—are systematically left unprotected. This is precisely backwards: CARG protects responses that need it least.

### v3

@@ 1 Introduction
  • Confidence calibration failure. We show that confidence poorly predicts correctness in reasoning models ($r=0.07$, $p=0.079$; ROC-AUC = 0.57), with systematic overconfidence induced by extended reasoning traces undermining confidence-based interventions.
@@ 4.1 Reasoning Models Are More Robust
Results. As shown in Table 2, all reasoning models outperform the GPT-4o baseline on $Acc_{\text{init}}$ (82–95% vs. 78%), confirming stronger baseline factual knowledge. For multi-turn consistency, the majority of reasoning models show substantial improvements ($Acc_{\text{avg}}$: 95–99% vs. 91.3%; PWC: 1.746–1.797 vs. 1.693), with several exhibiting $Acc_{\text{avg}}$ exceeding $Acc_{\text{init}}$, suggesting they leverage re-reasoning opportunities for error recovery.
@@
GPT-5.2 Reasoning 82.29% 96.31% 1.766
GPT-5.1 Reasoning 82.57% 98.92% 1.780
GPT-OSS Reasoning 88.71% 98.53% 1.795
Grok-4.1 Reasoning 92.43% 97.06% 1.797
DeepSeek-R1 Reasoning 91.86% 89.91% 1.727
Claude-4.5 Reasoning 94.86% 86.31% 1.675
Grok-3 Reasoning 85.29% 97.72% 1.772
Gemini-2.5 Reasoning 91.43% 96.48% 1.755
Qwen-3 Reasoning 89.86% 95.01% 1.746
GPT-4o Baseline 78.14% 91.29% 1.693
@@
We conducted Welch’s $t$-tests comparing each reasoning model’s per-question PWC scores against the GPT-4o baseline. Eight of nine reasoning models demonstrated significantly higher PWC than the baseline: GPT-OSS ($t=6.38$, $p<0.001$, $d=0.38$), Grok-4.1 ($t=6.60$, $p<0.001$, $d=0.40$), GPT-5.1 ($t=5.21$, $p<0.001$, $d=0.31$), GPT-5.2 ($t=4.23$, $p<0.001$, $d=0.25$), Grok-3 ($t=4.49$, $p<0.001$, $d=0.27$), Gemini-2.5 ($t=3.59$, $p<0.001$, $d=0.21$), Qwen-3 ($t=2.94$, $p=0.003$, $d=0.17$), and DeepSeek-R1 ($t=2.00$, $p=0.046$, $d=0.12$). An aggregate one-sample $t$-test confirmed that reasoning models as a class exhibit significantly higher PWC than the baseline (mean $\Delta=+0.064$, $t=5.02$, $p=0.001$).
@@ Results.
The confidence-correctness relationship is weak. The point-biserial correlation between per-question average confidence and correctness is $r=0.07$ ($p=0.079$), failing to reach significance at $\alpha=0.05$. Using confidence as a classifier for correctness yields ROC-AUC of 0.57, barely above chance. The confidence distribution itself reveals the problem: scores cluster tightly with mean 93.5%, standard deviation 4.4%, and range 75% to 100%.
@@
Low (<<33rd percentile) 9.3% 193
Medium (33rd–67th) 6.7% 193
High (>>67th percentile) 6.2% 193
@@
Large reasoning models exhibit systematic overconfidence: confidence scores cluster tightly around 93–95% regardless of actual correctness, yielding a compressed distribution with poor discriminative power. The model is nearly as confident about incorrect answers as correct ones.
Table 5 reveals a critical problem: low-confidence correct answers flip at the highest rate (9.3%), yet these are precisely the responses that confidence-based interventions like CARG would not protect. This creates a systematic selection bias—CARG protects already-robust high-confidence responses while leaving the most vulnerable responses exposed.
@@
Confidence is a poor proxy for correctness in large reasoning models ($r=0.07$, n.s.), undermining the core assumption of CARG-style interventions.
@@
Method $Acc_{\text{avg}}$ PWC
Baseline (No CARG) 98.50% 1.765
overall CARG 98.39% 1.773
answer_only CARG 98.29% 1.769
random CARG 98.88% 1.778
Table 6: CARG performance by confidence elicitation method (GPT-5.1). PWC uses the same exponential discounting formula as Table 2. Counterintuitively, random outperforms structured extraction approaches.
@@
(1) Overconfidence undermines targeted selection. With confidence scores clustering at 93–95% for both overall and answer_only, neither method provides meaningful discrimination. The signal-to-noise ratio is too low: most responses fall above any reasonable threshold, negating the selective benefit of confidence-based intervention.
(2) Selection bias amplifies vulnerability. Structured CARG preferentially protects high-confidence responses, which are already relatively robust (6.2% flip rate). Meanwhile, low-confidence correct answers—the most vulnerable group (9.3% flip rate)—are systematically left unprotected. This is precisely backwards: CARG protects responses that need it least.